# DRAGON: Distributional Rewards Optimize Diffusion Generative Models

### v1

@@ 5.6 Human Listening Test
We perform a listening test for subjective evaluation using two DRAGON models from Table 1 – one optimizing aesthetics score and the other optimizing per-song VAE-FAD to ALIM audio. We instruct 21 raters to compare the overall quality of blinded music pairs. Each rater is given 40 independently selected random SDNV prompts, along with the corresponding generation pairs. Out of each pair, one piece is from our baseline model and the other is from a DRAGON model (20 pairs for each DRAGON model). Clips are loudness-normalized to $-23\textrm{dB}$ LUFS, and presented in random order. To accelerate the test, clips are randomly cropped into 5-second snippets (same start/end timestamp per pair).
Across all raters, both DRAGON models outperform the baseline, with DRAGON-aesthetics achieving a 60.2% human-labeled win rate and DRAGON-VAE-FAD managing 61.0%. Despite DRAGON-aesthetics receiving a higher machine-predicted aesthetics win rate (85.2%) than DRAGON-VAE-FAD (78.3%), its human-perceived quality is slightly worse. This is likely because DRAGON-aesthetics incurs some overfitting by directly maximizing the predicted aesthetics. In contrast, DRAGON-VAE-FAD learns in an instance- to-distribution approach, reducing overfitting by not relying on the DMA preference dataset. In summary, DRAGON improves human-perceived music quality with sparse human feedback (via our aesthetics model) and even with no human feedback (via per-song FAD). In Section B.4, we use statistical hypothesis testing to confirm our improvement and derive a 95% confidence win rate lower bound.
@@ A.1 Diffusion-DPO and Diffusion-KTO Loss
Specifically, consider the forward diffusion process, where we inject Gaussian noise into a demonstration $x_{0}$ to form a noisy example $x_{t}$. That is, we randomly select a diffusion time step $t$ from a fixed distribution ${\mathcal{T}}$ supported on $[0,t_{\max}]$. Then, we sample $x_{t}$ from a Gaussian distribution $q(x_{t}|x_{0})={\mathcal{N}}(x_{t};\alpha_{t}x_{0},\sigma_{t}^{2}I)$, where $\alpha_{t}\in[0,1]$ and $\sigma_{t}$ represent the noise scheduling (Ho et al., 2020; Song et al., 2021a; Karras et al., 2022). Next, we query our model $f_{\theta}$ to denoise from $x_{t}$ and obtain the result $f_{\theta}(x_{t},t)$, where we omit the conditional information (prompt) in the notation for simplicity. Wallace et al. (2024) showed $\log\pi_{\theta}(x_{0})$ can be maximized via minimizing the surrogate objective function
@@
Practical algorithms that optimize the above quantity often incorporate weighting and regularization terms to stabilize training (Yang et al., 2024; Wallace et al., 2024; Li et al., 2024; Hong et al., 2024). Among these algorithms, Diffusion-DPO and Diffusion-KTO are two of the most popular examples. Diffusion-DPO, which requires paired demonstrations, solves
$\max_{\theta}\;\ \mathbb{E}_{\begin{subarray}{c}(x_{+0},x_{-0})\sim({\mathcal{D}}_{+},{\mathcal{D}}_{-}),\;\ t\sim{\mathcal{T}}\\ x_{+t}\sim q(x_{+t}|x_{+0}),\;\ x_{-t}\sim q(x_{-t}|x_{-0})\end{subarray}}\Big{[}\sigma\Big{(}\beta(t)\cdot\big{(}A_{\theta}(x_{+0},x_{+t},t)-A_{\theta}(x_{-0},x_{-t},t)\big{)}\Big{)}\Big{]},\vspace{-5mm}$ (6)
@@
Here, $A_{\theta}(x_{0},x_{t},t)$ represents how much $f_{\theta}$’s de-noising result is closer to the noiseless demonstration than $f_{\textrm{ref}}$’s result, where $f_{\textrm{ref}}$ is a “reference model” used for regularization. In practice, the pre-trained base model before DRAGON fine-tuning is used as $f_{\textrm{ref}}$. Intuitively, eq. 6 pulls the de-noising results from $f_{\theta}$ towards corresponding demonstrations in ${\mathcal{D}}_{+}$ and pushes them away from examples in ${\mathcal{D}}_{-}$.
@@
where the distribution ${\mathcal{D}}_{r}$ is randomly chosen between ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$, the binary variable $\mathrm{sgn}({\mathcal{D}}_{r}={\mathcal{D}}_{+})$ is $1$ when ${\mathcal{D}}_{r}$ is ${\mathcal{D}}_{+}$ and $-1$ when ${\mathcal{D}}_{r}$ is ${\mathcal{D}}_{-}$, the term $A_{\theta}(x_{0},x_{t},t)$ is defined in eq. 7, and $\bar{A}_{\theta}$ is a regularization term. Specifically, $\bar{A}_{\theta}$ is obtained by sampling an independent batch of $x_{0}^{(1)},\ldots,x_{0}^{(m)}\sim{\mathcal{D}}_{r}$, $t^{(1)},\ldots,t^{(m)}\sim{\mathcal{T}}$, and $x_{t}^{(1)},\ldots,x_{t}^{(m)}\sim q(x_{t}|x_{0}^{(1)}),\ldots,q(x_{t}|x_{0}^{(m)})$ and computing the “average $A_{\theta}(\cdot)$ value” via the formula
@@
In Section 5.2, we present ablation studies between Diffusion-KTO and Diffusion-DPO and between paired and unpaired demonstrations.
@@ A.2 Other Loss Functions and Beyond Binary ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$
Using DRAGON, we can incorporate other loss functions that learn from binary demonstrations without modification. Beyond Diffusion-DPO (Wallace et al., 2024) and Diffusion-KTO (Li et al., 2024), such loss functions include MaPO (Hong et al., 2024) and D3PO (Yang et al., 2024).
DRAGON can also scale beyond binary demonstrations. For reward functions like $r_{\textrm{instance}}$ that evaluate individual generations, this extension is straightforward. Loss functions for binary demonstrations generally decide the direction of optimization via the positive/negative nature of a demonstration. That is, they explicitly or implicitly involve the $\textrm{sgn}({\mathcal{D}}_{r}={\mathcal{D}}_{+})$ term introduced in the Diffusion-KTO loss eq. 8. Since ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$ are formed via a reward-based splitting operation, if we focus on a particular example $x_{0}$ sampled from the demonstrative distribution ${\mathcal{D}}_{r}$, then $\textrm{sgn}({\mathcal{D}}_{r}={\mathcal{D}}_{+})$ is equivalent to $\textrm{sgn}(r_{\textrm{instance}}(x_{0})>r_{\textrm{threshold}})$, where $r_{\textrm{threshold}}$ is the splitting threshold. We can extend beyond the binary preference assumption by replacing the discontinuous sign function with a continuous function, such as sigmoid or identity. One simple continuous-reward loss function is thus
@@
which is equivalent to the DPOK objective with the KL-D setting (Fan et al., 2024), making it an example of reward-weighted regression algorithms (Peters & Schaal, 2007). That is, DPOK can be regarded as a special case of DRAGON under the instance-level reward scenario with non-binary preferences. One can similarly “continuize” the Diffusion-KTO loss as
@@
Alternatively, one can consider other traditional RL loss functions for diffusion models like DDPO (Black et al., 2024) or derive variants of GRPO (Shao et al., 2024) specialized for diffusion models,
@@
It is also possible to scale the number of demonstrative distributions beyond two by modifying Algorithm 1.
@@ B.1 The DMA Music Preference Dataset
The collection pipeline of our DMA preference dataset is a two-phase process. In Phase 1, users interact with a collection of music generation models via an interface. After receiving the user prompt, the interface generates a piece, which the user rates on a scale of 1-5. In Phase 2, we reuse the user prompts from Phase 1 and generate additional music pieces. We provide four examples per prompt, which the user again rates on a scale of 1-5, providing a direct contrastive signal. To enhance data diversity, Phase 2 also randomly mixes in some LLM-created prompts and ALIM training set captions. During DRAGON fine-tuning, generation quality is expected to improve. To help our aesthetics model generalize and mitigate the likelihood for DRAGON to quickly become out of distribution, we additionally mix in some high-quality human-created music. Specifically, for ALIM prompts used in Phase 2, we randomly mix in ground-truth ALIM music.
@@
Our final dataset consists of 800 prompts, 1,676 music pieces with various durations totaling 15.97 hours, and 2,301 ratings from 63 raters (multiple raters can rate the same generation). The proportion of each prompt source is shown in Table 2. Due to the small dataset size, we do not explicitly disentangle different aspects of music quality and text correspondence, and instead ask for a single overall opinion rating. Despite our dataset being orders of magnitude smaller than comparable modern image aesthetics datasets (see Table 3 for detailed comparisons), we will show that by leveraging DRAGON’s versatile and on-policy training pipeline, we can improve human-perceived generation quality with high data efficiency.
@@
SAC (Pressman et al., 2022) Image >238,000 AI-generated Human-rated 1-10 score
AVA (Murray et al., 2012) Image >250,000 Human-created Human-rated 1-10 score
LAION-Aes V2 (Schuhmann, 2022) Image 1.2 Billion Human-created Model-predicted 1-10 score
Pick-a-Pic (Kirstain et al., 2023) Image >1 Million AI-generated Human-rated Paired binary
RichHF-18K (Liang et al., 2024) Image 18,000 AI-generated Human-rated 1-5 multi-facet
BATON (Liao et al., 2024) In-the-wild Audio 2,763 AI-generated Human-rated Paired binary
Audio-Alpaca (Majumder et al., 2024) In-the-wild Audio 15,000 AI-generated Model-predicted Paired binary
MusicRL (Cideron et al., 2024) Music 285,000 AI-generated Human-rated Paired binary
Ours Music 1,676 Mostly AI-generated Human-rated 1-5 score
@@ B.2 Aesthetics Predictor Details and Ablations
To verify the performance of the aesthetics predictor and perform ablation studies, we split the DMA dataset into train/validation subsets by an 85/15 ratio. Specifically, we use overall PLCC and per-prompt SRCC to quantify the agreement between predicted aesthetics and human ratings on the validation split. PLCC is a number between $-1$ and $1$ that represents the “normalized covariance” (a larger value means more positive correlation and zero implies no correlation). SRCC is defined as the PLCC over rankings, and “per-prompt” means computing the SRCC among all generations from each prompt and averaging across all prompts. Per-prompt SRCC more accurately analyzes reward signals’ agreement when comparing generations from the same prompt. Since we learn from contrastive demonstrations ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$ evaluated with reward signals (often paired by prompt), per-prompt SRCC is particularly meaningful, as a higher per-prompt SRCC makes the aesthetics predictor more reliable as a reward function.
Ablation 1: Embedding averaging. As mentioned in Section 5.1, there are numerous ways to pre-process our 32-second 44.1kHz generations into the 10-second 48kHz format required by CLAP. We start with the original CLAP inference setting introduced in (Wu* et al., 2023) and explore several modifications:
@@
As shown in Figure 5, all settings obtain similar overall PLCC, with Setting 4 achieving the best per-prompt SRCC. Recall that the overall PLCC is computed across all music from all captions, entangling semantics and aesthetics, whereas per-prompt SRCC removes the influence of the semantic context by unifying the prompt. The low-SRCC settings tend to pay more attention to the semantics information dictated by the prompt and infer the aesthetics score via semantics-aesthetics correlation. In contrast, high-SRCC settings directly learn aesthetics information, making their outputs more closely aligned with human preference when ranking generations from the same prompt (which is the aesthetics model’s task in DRAGON). In conclusion, while the regression quality is similar among all settings, different training/inference settings make the models learn different information. Hence, when training regressive aesthetics models, we need to look beyond typical full-dataset metrics and evaluate in settings that mirror a given use case. In our case, Setting 4 is the most reliable online music evaluation protocol.
Figure 5 also shows that the rate-then-average paradigm produces noticeably better results than CLAP’s original average-then-rate approach proposed in (Wu* et al., 2023). This makes sense because when averaged across all pieces in the DMA dataset, the within-piece standard deviation among Setting 4’s nine raw aesthetics predictions is $0.238$. For reference, the dataset-wide aesthetics prediction standard deviation is $0.755$, meaning that the within-piece variance is non-trivial. That is, different chunks of the same music piece can vary noticeably in predicted aesthetics score, and the rate-then-average approach mitigates this variance via ensembling, resulting in better performance.
@@
    Setting 4 with $[-0.5,0.5]$ peak normalization. Use the same training and inference setting as Setting 4, except we apply peak normalization to each 10-second chunk before encoding them, so that all waveforms are in $[-0.5,0.5]$.
@@
    Setting 4 with $[-1,1]$ peak normalization. Same as Setting 4b, except the peak normalization scales each 10-second chunk to $[-1,1]$ instead of $[-0.5,0.5]$.
As shown in Figure 5, Setting 4b ($[-0.5,0.5]$) achieves a much higher per-prompt SRCC than Setting 4c ($[-1,1]$). While Setting 4b’s per-prompt SRCC is slightly lower than Setting 4, we believe the magnitude-invariance guarantee outweighs the minor performance drop, and thus keep the peak normalization. Note that the diffusion VAE encoder, which doubles as a mapping to the latent diffusion space and an FAD reference embedding extractor, does not use peak normalization to preserve information for audio reconstruction.
Ablation 3: Label normalization. We perform label normalization to transform our 1-5 raw human ratings into a well-conditioned zero-mean distribution and study two settings:
@@
    Settings 4-G and 4b-G – Per-rater normalization with Gibbs sampler. Drawing inspiration from Hierarchical Bayesian Model (Lindley & Smith, 1972), we assume that the “rater harshness”, quantified by a rater’s average rating, forms a standard Gaussian distribution. The ratings from each rater then form another Gaussian distribution centered around this average. To effectively estimate the “average harshness” and the rating variance of each rater, we utilize the Gibbs sampler (Gelfand et al., 1990). Via rater harshness modeling, we can perform a more fine-grained per-rater normalization.
As shown in Figure 5, per-rater label normalization does not always outperform conventional global normalization, with 4b-G better than 4b but Setting 4 superior to Setting 4-G. While per-rater normalization is more sophisticated, a time-invariant Gaussian distribution may not accurately model each rater’s ratings due to small sample size and potential preference change over time. As a result, global normalization, with the less precise overall Gaussian modeling, remains competitive.
Our final selection is Setting 4b, which reliably generalizes to unseen data while being simple. On the validation split of the DMA dataset, Setting 4b achieves an overall PLCC of $0.576$ and a per-prompt SRCC of $0.484$. For DRAGON training and evaluation, we remove the train/validation split and use Setting 4b to train a new model with all available data. Figure 6 presents this new model’s prediction distribution over the DMA dataset.
Comparison with other aesthetics models. With Setting 4b, the performance of our preference model is on par with existing preference models of other modalities trained with much larger data sizes. For example, Liang et al. (2024) trained a sophisticated multi-branch image rating model that simultaneously predicts aesthetics score and several other metrics. With a similar 1-5 scoring scale, this model is one of the most comparable works to ours. Trained on about 16,000 rated images, the aesthetics prediction in (Liang et al., 2024) achieves a PLCC of $0.605$. Despite our DMA preference dataset containing one order of magnitude fewer ratings, our model achieves a similar validation set PLCC of $0.576$.

### v2

@@ 5.6 Human Listening Test
We perform a listening test for subjective evaluation using two DRAGON models from Table˜2 – one optimizing aesthetics score and the other optimizing per-song VAE-FAD to ALIM audio. We instruct 21 raters to compare the overall quality of blinded music pairs. Each rater is given 40 independently selected random SDNV prompts, along with the corresponding generation pairs. Out of each pair, one piece is from our baseline model and the other is from a DRAGON model (20 pairs for each DRAGON model). Clips are loudness-normalized to $23\textrm{dB}$ LUFS, and presented in random order. To accelerate the test, clips are randomly cropped into 5-second snippets (same start/end timestamp per pair).
Across all raters, both DRAGON models outperform the baseline, with DRAGON-aesthetics achieving a 60.2% human-labeled win rate and DRAGON-VAE-FAD managing 61.0%. Despite DRAGON-aesthetics receiving a higher machine-predicted aesthetics win rate (85.2%) than DRAGON-VAE-FAD (78.3%), its human-perceived quality is slightly worse. This is likely because DRAGON-aesthetics incurs some overfitting by directly maximizing the predicted aesthetics. In contrast, DRAGON-VAE-FAD learns in an instance- to-distribution approach, reducing overfitting by not relying on the DMA preference dataset. In summary, DRAGON improves human-perceived music quality with sparse human feedback (via our aesthetics model) and even with no human feedback (via per-song FAD). In Section˜B.4, we use statistical hypothesis testing to confirm our improvement and derive a 95% confidence win rate lower bound.
@@ A.1 Diffusion-DPO and Diffusion-KTO Loss
Specifically, consider the forward diffusion process, where we inject Gaussian noise into a demonstration $x_{0}$ to form a noisy example $x_{t}$. That is, we randomly select a diffusion time step $t$ from a fixed distribution $\mathcal{T}$} supported on $0,t_{\max}]$. Then, we sample $x_{t}$ from a Gaussian distribution $q(x_{t}|x_{0})={\mathcal{N}}(x_{t};\alpha_{t}x_{0},\sigma_{t}^{2}I)$, where $\alpha_{t}\in[0,1]$ and $\sigma_{t}$ represent the noise scheduling (DDPM; DDIM; EDM). Next, we query our model $f_{\theta}$ to denoise from $x_{t}$ and obtain the result $f_{\theta}(x_{t},t)$, where we omit the conditional information (prompt) in the notation for simplicity. Diffusion-DPO showed $\log\pi_{\theta}(x_{0})$ can be maximized via minimizing the surrogate objective function
@@
Practical algorithms that optimize the above quantity often incorporate weighting and regularization terms to stabilize training (D3PO; Diffusion-DPO; Diffusion-KTO; MaPO). Among these algorithms, Diffusion-DPO and Diffusion-KTO are two of the most popular examples. Diffusion-DPO, which requires paired demonstrations, solves
$\max_{\theta}\;\ \mathbb{E}_{\begin{subarray}{c}(x_{+0},x_{-0})\sim({\mathcal{D}}_{+},{\mathcal{D}}_{-}),\;\ t\sim{\mathcal{T}}\\ x_{+t}\sim q(x_{+t}|x_{+0}),\;\ x_{-t}\sim q(x_{-t}|x_{-0})\end{subarray}}\Big[\sigma\Big(\beta(t)\cdot\big(A_{\theta}(x_{+0},x_{+t},t)-A_{\theta}(x_{-0},x_{-t},t)\big)\Big)\Big],\vskip-14.22636pt$ (6)
@@
Here, $A_{\theta}(x_{0},x_{t},t)$ represents how much $f_{\theta}$’s de-noising result is closer to the noiseless demonstration than $f_{\textrm{ref}}$’s result, where $f_{\textrm{ref}}$ is a “reference model” used for regularization. In practice, the pre-trained base model before DRAGON fine-tuning is used as $f_{\textrm{ref}}$. Intuitively, eq.˜6 pulls the de-noising results from $f_{\theta}$ towards corresponding demonstrations in $\mathcal{D}$$_{+}$ and pushes them away from examples in $\mathcal{D}$$_{-}$.
@@
where the distribution ${\mathcal{D}}_{r}$ is randomly chosen between $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$, the binary variable $\mathrm{sgn}({\mathcal{D}}_{r}={\mathcal{D}}_{+})$ is $1$ when ${\mathcal{D}}_{r}$ is $\mathcal{D}$$_{+}$ and 1-1 when ${\mathcal{D}}_{r}$ is $\mathcal{D}$$_{-}$, the term $A_{\theta}(x_{0},x_{t},t)$ is defined in eq.˜7, and $\bar{A}_{\theta}$ is a regularization term. Specifically, $\bar{A}_{\theta}$ is obtained by sampling an independent batch of $x_{0}^{(1)},\ldots,x_{0}^{(m)}\sim{\mathcal{D}}_{r}$, $t^{(1)},\ldots,t^{(m)}\sim{\mathcal{T}}$, and $x_{t}^{(1)},\ldots,x_{t}^{(m)}\sim q(x_{t}|x_{0}^{(1)}),\ldots,q(x_{t}|x_{0}^{(m)})$ and computing the “average $A_{\theta}(\cdot)$ value” via the formula
@@
In Section˜5.2, we present ablation studies between Diffusion-KTO and Diffusion-DPO and between paired and unpaired demonstrations.
@@ A.2 Other Loss Functions and Beyond Binary $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$
Using DRAGON, we can incorporate other loss functions that learn from binary demonstrations without modification. Beyond Diffusion-DPO (Diffusion-DPO) and Diffusion-KTO (Diffusion-KTO), such loss functions include MaPO (MaPO) and D3PO (D3PO).
DRAGON can also scale beyond binary demonstrations. For reward functions like $r_{\textrm{instance}}$ that evaluate individual generations, this extension is straightforward. Loss functions for binary demonstrations generally decide the direction of optimization via the positive/negative nature of a demonstration. That is, they explicitly or implicitly involve the $\textrm{sgn}({\mathcal{D}}_{r}={\mathcal{D}}_{+})$ term introduced in the Diffusion-KTO loss eq.˜8. Since $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$ are formed via a reward-based splitting operation, if we focus on a particular example $x_{0}$ sampled from the demonstrative distribution ${\mathcal{D}}_{r}$, then $\textrm{sgn}({\mathcal{D}}_{r}={\mathcal{D}}_{+})$ is equivalent to $\textrm{sgn}(r_{\textrm{instance}}(x_{0})>r_{\textrm{threshold}})$, where $r_{\textrm{threshold}}$ is the splitting threshold. We can extend beyond the binary preference assumption by replacing the discontinuous sign function with a continuous function, such as sigmoid or identity. One simple continuous-reward loss function is thus
@@
which is equivalent to the DPOK objective with the KL-D setting (DPOK), making it an example of reward-weighted regression algorithms (RWR). That is, DPOK can be regarded as a special case of DRAGON under the instance-level reward scenario with non-binary preferences. One can similarly “continuize” the Diffusion-KTO loss as
@@
Alternatively, one can consider other traditional RL loss functions for diffusion models like DDPO (DDPO) or derive variants of GRPO (DeepSeekMath) specialized for diffusion models,
@@
It is also possible to scale the number of demonstrative distributions beyond two by modifying Algorithm˜1.
@@ B.1 The DMA Music Preference Dataset
The collection pipeline of our DMA preference dataset is a two-phase process. In Phase 1, users interact with a collection of music generation models via an interface. After receiving the user prompt, the interface generates a piece, which the user rates on a scale of 1–5. In Phase 2, we reuse the user prompts from Phase 1 and generate additional music pieces. We provide four examples per prompt, which the user again rates on a scale of 1–5, providing a direct contrastive signal. To enhance data diversity, Phase 2 also randomly mixes in some LLM-created prompts and ALIM training set captions. During DRAGON fine-tuning, generation quality is expected to improve. To help our aesthetics model generalize and mitigate the likelihood for DRAGON to quickly become out of distribution, we additionally mix in some high-quality human-created music. Specifically, for ALIM prompts used in Phase 2, we randomly mix in ground-truth ALIM music.
@@
Our final dataset consists of 800 prompts, 1,676 music pieces with various durations totaling 15.97 hours, and 2,301 ratings from 63 raters (multiple raters can rate the same generation). The proportion of each prompt source is shown in Table˜3. Due to the small dataset size, we do not explicitly disentangle different aspects of music quality and text correspondence, and instead ask for a single overall opinion rating. Despite our dataset being orders of magnitude smaller than comparable modern image aesthetics datasets (see Table˜4 for detailed comparisons), we will show that by leveraging DRAGON’s versatile and on-policy training pipeline, we can improve human-perceived generation quality with high data efficiency.
@@
SAC (SAC) Image >238,000 AI-generated Human-rated 1-10 score
AVA (AVA) Image >250,000 Human-created Human-rated 1-10 score
LAION-Aes V2 (LAION-Aesthetics-Dataset) Image 1.2 Billion Human-created Model-predicted 1-10 score
Pick-a-Pic (Pick-A-Pic) Image >1 Million AI-generated Human-rated Paired binary
RichHF-18K (liang2024rich) Image 18,000 AI-generated Human-rated 1–5 multi-facet
BATON (BATON) In-the-wild Audio 2,763 AI-generated Human-rated Paired binary
Audio-Alpaca (TANGO2) In-the-wild Audio 15,000 AI-generated Model-predicted Paired binary
MusicRL (MusicRL) Music 285,000 AI-generated Human-rated Paired binary
Ours Music 1,676 Mostly AI-generated Human-rated 1–5 score
@@ B.2 Aesthetics Predictor Details and Ablations
To verify the performance of the aesthetics predictor and perform ablation studies, we split the DMA dataset into train/validation subsets by an 85/15 ratio. Specifically, we use overall PLCC and per-prompt SRCC to quantify the agreement between predicted aesthetics and human ratings on the validation split. PLCC is a number between 1-1 and $1$ that represents the “normalized covariance” (a larger value means more positive correlation and zero implies no correlation). SRCC is defined as the PLCC over rankings, and “per-prompt” means computing the SRCC among all generations from each prompt and averaging across all prompts. Per-prompt SRCC more accurately analyzes reward signals’ agreement when comparing generations from the same prompt. Since we learn from contrastive demonstrations $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$ evaluated with reward signals (often paired by prompt), per-prompt SRCC is particularly meaningful, as a higher per-prompt SRCC makes the aesthetics predictor more reliable as a reward function.
Ablation 1: Embedding averaging. As mentioned in Section˜5.1, there are numerous ways to pre-process our 32-second 44.1kHz generations into the 10-second 48kHz format required by CLAP. We start with the original CLAP inference setting introduced in (laionclap2023) and explore several modifications:
@@
As shown in Figure˜5, all settings obtain similar overall PLCC, with Setting 4 achieving the best per-prompt SRCC. Recall that the overall PLCC is computed across all music from all captions, entangling semantics and aesthetics, whereas per-prompt SRCC removes the influence of the semantic context by unifying the prompt. The low-SRCC settings tend to pay more attention to the semantics information dictated by the prompt and infer the aesthetics score via semantics-aesthetics correlation. In contrast, high-SRCC settings directly learn aesthetics information, making their outputs more closely aligned with human preference when ranking generations from the same prompt (which is the aesthetics model’s task in DRAGON). In conclusion, while the regression quality is similar among all settings, different training/inference settings make the models learn different information. Hence, when training regressive aesthetics models, we need to look beyond typical full-dataset metrics and evaluate in settings that mirror a given use case. In our case, Setting 4 is the most reliable online music evaluation protocol.
Figure˜5 also shows that the rate-then-average paradigm produces noticeably better results than CLAP’s original average-then-rate approach proposed in (laionclap2023). This makes sense because when averaged across all pieces in the DMA dataset, the within-piece standard deviation among Setting 4’s nine raw aesthetics predictions is $0.238$. For reference, the dataset-wide aesthetics prediction standard deviation is $0.755$, meaning that the within-piece variance is non-trivial. That is, different chunks of the same music piece can vary noticeably in predicted aesthetics score, and the rate-then-average approach mitigates this variance via ensembling, resulting in better performance.
@@
    Setting 4 with [0.5,0.5][-0.5,0.5] peak normalization. Use the same training and inference setting as Setting 4, except we apply peak normalization to each 10-second chunk before encoding them, so that all waveforms are in [0.5,0.5][-0.5,0.5].
@@
    Setting 4 with [1,1][-1,1] peak normalization. Same as Setting 4b, except the peak normalization scales each 10-second chunk to [1,1][-1,1] instead of [0.5,0.5][-0.5,0.5].
As shown in Figure˜5, Setting 4b ([0.5,0.5][-0.5,0.5]) achieves a much higher per-prompt SRCC than Setting 4c ([1,1][-1,1]). While Setting 4b’s per-prompt SRCC is slightly lower than Setting 4, we believe the magnitude-invariance guarantee outweighs the minor performance drop, and thus keep the peak normalization. Note that the diffusion VAE encoder, which doubles as a mapping to the latent diffusion space and an FAD reference embedding extractor, does not use peak normalization to preserve information for audio reconstruction.
Ablation 3: Label normalization. We perform label normalization to transform our 1–5 raw human ratings into a well-conditioned zero-mean distribution and study two settings:
@@
    Settings 4-G and 4b-G – Per-rater normalization with Gibbs sampler. Drawing inspiration from Hierarchical Bayesian Model (HBM), we assume that the “rater harshness”, quantified by a rater’s average rating, forms a standard Gaussian distribution. The ratings from each rater then form another Gaussian distribution centered around this average. To effectively estimate the “average harshness” and the rating variance of each rater, we utilize the Gibbs sampler (Gibbs). Via rater harshness modeling, we can perform a more fine-grained per-rater normalization.
As shown in Figure˜5, per-rater label normalization does not always outperform conventional global normalization, with 4b-G better than 4b but Setting 4 superior to Setting 4-G. While per-rater normalization is more sophisticated, a time-invariant Gaussian distribution may not accurately model each rater’s ratings due to small sample size and potential preference change over time. As a result, global normalization, with the less precise overall Gaussian modeling, remains competitive.
Our final selection is Setting 4b, which reliably generalizes to unseen data while being simple. On the validation split of the DMA dataset, Setting 4b achieves an overall PLCC of $0.576$ and a per-prompt SRCC of $0.484$. For DRAGON training and evaluation, we remove the train/validation split and use Setting 4b to train a new model with all available data. Figure˜6 presents this new model’s prediction distribution over the DMA dataset.
Comparison with other aesthetics models. With Setting 4b, the performance of our preference model is on par with existing preference models of other modalities trained with much larger data sizes. For example, liang2024rich trained a sophisticated multi-branch image rating model that simultaneously predicts aesthetics score and several other metrics. With a similar 1–5 scoring scale, this model is one of the most comparable works to ours. Trained on about 16,000 rated images, the aesthetics prediction in (liang2024rich) achieves a PLCC of $0.605$. Despite our DMA preference dataset containing one order of magnitude fewer ratings, our model achieves a similar validation set PLCC of $0.576$.